OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <8, 7>, v = <9, 7>

OpenStudy (zehanz):

You could use the formula:\[\cos \theta=\frac{ \vec u \cdot \vec v}{ |\vec u||\vec v| }\]assuming you are familiar with the dot product of vectors.

OpenStudy (anonymous):

here it is on mathematica.

OpenStudy (anonymous):

that wouldnt let me open it @Sshmoo

OpenStudy (anonymous):

im just learning it so im still so comfused

OpenStudy (anonymous):

i'll try converting it to a different file, all it is is a graph of the vectors and a dot product. It just looks shinier.

OpenStudy (anonymous):

ohok

OpenStudy (anonymous):

here it is in PDF

OpenStudy (anonymous):

okk hm so

OpenStudy (anonymous):

so which of these would the answer be ? -8.3° 1.7° 3.3° 13.3°

OpenStudy (zehanz):

If you calculate cos theta, you get: \[\cos \theta=\frac{ 8 \cdot 9 + 7 \cdot 7 }{ \sqrt{8^2+7^2}\sqrt{9^2+7^2} }=\frac{ 72+49 }{\sqrt{113}\sqrt{130} }=\frac{ 121 }{ \sqrt{113} \sqrt{130}}\approx 0.99833\]Now take the inverse cosine (cos^-1 on your calculator) to see the answer.

OpenStudy (anonymous):

THANK YOU!

OpenStudy (zehanz):

yw!